Determine the missing amount for each of the following:

Assets = Liabilities + Stockholders Equity

A. X = 715000 + 1885000

B. 600000 = X + 10000

C. 112400 = 23750 + X
Answer #1

A. X = 715,000+1,885,000

=$2,600,000

B.X= 600,000-10,000=$590,000

C.X=112400-23750=$88650.